Find the pH of the following :

0.50 M solution of a base with a Kb = 1.75 x 10-4
Answer #1

According to the condition given in let us write the reaction. - A + HQO - HAT OH INITIAL 0,5M EQUILIBRIUM 0.5-x x xKb = [HAJ[OH-] 1.75X 10% = CA] x? 0.5-X, 2 = [OH-] = 9.267x103M pots - log Low-] pot er log (9.267x103)POH = 2.03 pH + Pot = 14 PH = 14-2.03 PH - 11,97]